OpenStudy (anonymous):

If f(x) = 5x + 3, then what is f–1(x) ?

OpenStudy (bahrom7893):

\[y=5x+3\] Solve for x.

OpenStudy (bahrom7893):

What do you get?

OpenStudy (bahrom7893):

Hello?

OpenStudy (anonymous):

-3/5?

OpenStudy (campbell_st):

start with \[x = 5y + 3\] to find the inverse make y the subject

OpenStudy (anonymous):

5y=3-x

OpenStudy (campbell_st):

not quite x - 3 = 5y now divide both sides by 5

OpenStudy (anonymous):

1x/5-3/5=y

OpenStudy (campbell_st):

thats correct... or y = (x -3)/5

OpenStudy (anonymous):

so is that the answer?

OpenStudy (campbell_st):

yep... thats the inverse...

OpenStudy (anonymous):

ok. thanks can you hall me with a word problem

OpenStudy (campbell_st):

if you post it in the normal way I'll look at it... there are lots of people who can help

OpenStudy (anonymous):

ok